6x-5+8x-6=8x+6x+11
solve for x
please I am just begining algebra in college, and I am having a hard time

Answer by Alan3354(21583)

6x-5+8x-6=8x+6x+11
----------
Collect all the x's on both sides of the =
14x-5-6 = 14x+11
14x-11 = 14x + 11
That can't be, so there's no solution. No value of x that works (maybe infinity, or undefined).
It's sort of a trick question.